# MSFMamba: Multi-Scale Feature Fusion State Space Model for Multi-Source Remote Sensing Image Classification

### v1

@@ III-C Multi-Scale Spatial Mamba Block (MSpa-Mamba)
Multi-Scale Spatial SSM (MSpa-SSM). As shown in Fig. 1(b), MSpa-SSM is the critical part of the MSpa-Mamba block. In MSpa-SSM, features at multiple scales are generated through depth-wise convolution (DWConv) with different stride values. These multi-scale feature maps are then handled by four scanning routes within SSM [46]. The scanning routes are divided into two groups: Two maintain the original resolution and are processed by the SSM. The others are downsampled, processed by the SSM, and then upsampled. To be more specific, we use DWConv with strides of 1 and 2 to obtain feature map $\mathbf{Z}_{1}\in\mathbb{R}^{H\times W\times C}$ and $\mathbf{Z}_{2}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times C}$, respectively. Next, $\mathbf{Z}_{1}$ and $\mathbf{Z_{2}}$ are fed into the SSM as follows:
$[\mathbf{Y}_{1},\mathbf{Y}_{2}]=\textrm{SSM}(\sigma_{1}(\mathbf{Z}_{1}),\sigma_{1}(\mathbf{Z}_{2})),$ (9)
$[\mathbf{Y}_{3},\mathbf{Y}_{4}]=\textrm{SSM}(\sigma_{3}(\mathbf{Z}_{3}),\sigma_{4}(\mathbf{Z}_{4})),$ (10)
where $\sigma$ is the transformation that converts 2D feature maps into 1D sequences in SSM. $\mathbf{Y}$ is the processed sequence. The obtained sequences are converted back into 2D feature maps, and the downsampled feature maps are interpolated for merging as follows:
@@
where $\beta$ is the inverse transformation of $\sigma$. $\mathbf{Z}^{\prime}$ is the feature map enhanced by SSM. ’Inter’ represents the interpolation operation.
III-D Spectral Mamba Block (Spe-Mamba)
HSI covers a much larger spectral range and higher spectral resolution. How to model the relationship between spectra is an important problem. In this paper, we design a Spectral Mamba (Spe-Mamba) block for spectral feature modeling.
@@ III-D Spectral Mamba Block (Spe-Mamba)
It should be noted that the Spe-Mamba has similar structures with the MSpa-Mamba. The difference between Spe-Mamba and MSpa-Mamba is the SSM layer. In MSpa-SSM, the features are converted into 1D sequences along the spatial dimension. While in Spe-Mamba, the features are converted into 1D sequences along the channel dimension, and the features are only scanned in two directions.
@@
III-E Fusion Mamba Block (Fus-Mamba)
Traditional SSMs only handle single inputs, and they hardly combine the cross-modal information from multi-source data. To solve the problem, we propose Fusion Mamba (Fus-Mamba) block to integrate features from different modalities. As shown in Fig. 2, the HSI and LiDAR/SAR features undergo the same processing procedure, which is similar to the MSpa-Mamba and Spe-Mamba. Multi-source features are fed into the Fusion-SSM (Fus-SSM) block for cross-modal feature fusion and interaction.
The Fus-SSM is the critical part of Fus-Mamba, and details of the Fus-SSM are illustrated in Fig. 2(b). Fus-SSM has a symmetric structure. On the top part, sequenced HSI feature $\mathbf{F}_{h}\in\mathbb{R}^{P\times C}$ is responsible for generating the projection and timescale parameters, while the sequenced LiDAR/SAR feature $\mathbf{F}_{x}\in\mathbb{R}^{P\times C}$ is to be handled to generate the output feature $\mathbf{F}_{ho}$. The bottom part of Fus-SSM has a similar structure with the top part. $\mathbf{F}_{x}$ is responsible for generating the projection and timescale parameters, while $\mathbf{F}_{h}$ is to be processed to generate the output feature $\mathbf{F}_{xo}$.
Algorithm 1 shows the detailed computation process of $\mathbf{F}_{ho}$. A and D are initialized at the beginning, while B, C, and $\Delta$ are obtained through linear layers from the input sequence $\textbf{F}_{h}$. In the original Mamba, the sequence fed into SSM is input sequence $\textbf{F}_{h}$, while in the proposed Fus-SSM, the sequence fed into SSM is from another modality $\textbf{F}_{x}$. Since the computation of $\mathbf{F}_{xo}$ is very similar to that of $\mathbf{F}_{ho}$, detailed description will not be repeated here.
@@ III-E Fusion Mamba Block (Fus-Mamba)
5: /* A and D are randomly initialized. */
6:$\mathbf{B}:{\rm{{\color[rgb]{0,0.390625,0}(P,N)}}}\leftarrow\mathbf{Linear}(\mathbf{F}_{h})$
7:$\mathbf{C}:{\rm{{\color[rgb]{0,0.390625,0}(P,N)}}}\leftarrow\mathbf{Linear}(\mathbf{F}_{h})$
@@
12:$\mathbf{F}_{ho}\leftarrow{\rm{SSM}(\mathbf{\overline{A}},\mathbf{\overline{B}},\mathbf{C},\mathbf{D})}(\mathbf{F}_{x})$
@@ IV-A Dataset Description
To assess the effectiveness of our proposed MSFMamba, we employ it to three multi-source remote sensing datasets: Berlin, Augsburg, and Houston 2018. The Berlin and Augsburg datasets are utilized for hyperspectral and SAR data classification, while the Houston 2018 dataset is employed for hyperspectral and LiDAR data classification.
TABLE I: The number of training and test samples for the Berlin, Augsburg, and Houston2018 datasets.
@@ 20 Stadium seats 1000 26296
Berlin Dataset. It covers the urban and rural areas of Berlin. The hyperspectral data is simulated EnMAP data based on HyMap data. The SAR data is Sentinel-1 dual-Pol (VV–VH) single look complex (SLC) product obtained from the European Space Agency (ESA) [52]. The HSI has 797$\times$220 pixels, 244 spectral bands in the wavelength range of 400–2500 nm [53]. The SAR image has 1723$\times$476 pixels. The nearest neighbor interpolation was employed to match the spatial resolution of the HSI and SAR data.
@@
Table LABEL:table_dataset enumerates the number of samples for both training and testing on three datasets. Fig. 3 displays the HSI data in false color, along with the LiDAR/SAR images and the ground truth. The evaluation metrics used in this paper include Overall Accuracy (OA), Average Accuracy (AA), and Kappa. Specifically, OA denotes the proportion of accurately classified samples throughout the entirety of the dataset, AA signifies the mean accuracy across all classes present within the dataset, and Kappa serves as a statistical measure of the concurrence between predicted and true labels.
Figure 3: Visualization of the three datasets used in our experiments. (a) Berlin dataset. (b) Augsburg dataset. (c) Houston 2018 dataset.
@@ IV-B Parameter Analysis
Figure 7: Classification results of different methods for the Berlin dataset. (a) Ground Truth. (b) TBCNN. (c) FusAtNet. (d) $S^{2}$et. (e) DFINet. (f) AsyFFNet. (g) ExViT. (h) HCT. (i) MACN. (j) proposed MSFMamba.
Figure 8: Classification results of different methods for the Augsburg dataset. (a) Ground Truth. (b) TBCNN. (c) FusAtNet. (d) $S^{2}$et. (e) DFINet. (f) AsyFFNet. (g) ExViT. (h) HCT. (i) MACN. (j) Proposed MSFMamba.
Figure 9: Classification results of different methods for the Houston 2018 dataset. (a) Ground Truth. (b) TBCNN. (c) FusAtNet. (d) $S^{2}$et. (e) DFINet. (f) AsyFFNet. (g) ExViT. (h) HCT. (i) MACN. (j) Proposed MSFMamba.
@@
Number of Spatial-Spectral Mamba Module. The spatial-spectral Mamba module is the critical part in our MSFMamba, and the number of the module $L$ is a key parameter affecting the classification performance. We tested the optimal number of spatial-spectral Mamba module on each dataset. Fig. 4 reveals the relationship between the OA value and $L$ on each dataset. It shows that our MSFMamba achieves the best classification performance on the Augsburg dataset when $L=1$. When $L=2$, our MSFMamba achieves the best best classification performance on the Berlin and Houston2018 datasets. Therefore, in our following experiments, we use $L=1$ on the Augsburg dataset, and $L=2$ on the other datasets.
Number of Principal Components in HSI. We use PCA to reduce the spectral redundancy of the input HSI. The number of principal components $N_{p}$ is a critical parameter. We test different values of $N_{p}$ ranging from 20 to 40, and the experimental results are shown in Fig. 5. It can be seen that when $N_{p}=30$, the OA values for Augsburg and Houston2018 achieve the best, and when $N_{p}=35$, the OA value for Berlin achieves the best.
Number of States in SSM. The number of states is an important parameter in the state space model. We tested five different number of states ranging from 4 to 20. As shown in Fig. 6, the optimal number of states for the Houston2018 and Berlin datasets is 16, while for the Augsburg dataset, the optimal number of states is 8.
The best settings of these experiments were used to evaluate our MSFMamba for comparison with the other state-of-the-art methods. All experiments were performed with an RTX 4090 GPU and 32GB of RAM using PyTorch.
@@
Forest  79.51  86.24  76.22  76.90  81.02  62.71  78.56  77.44  81.05
Residential area  92.27  91.38  95.09  93.41  93.62  94.34  93.71  92.73  93.00
Industrial area  42.27  19.76  41.84  37.09  38.90  41.02  44.48  41.38  47.96
Low plants  72.10  20.00  68.53  62.76  71.52  68.79  72.79  74.56  75.81
Soil  65.05  48.72  60.19  66.99  71.08  60.62  74.73  77.29  79.16
Allotment  17.59  38.89  42.56  46.15  26.04  26.71  28.21  24.44  38.35
Commercial area  15.49  18.47  16.42  18.80  13.23  35.01  26.28  25.04  33.37
Water  66.37  29.61  68.00  66.94  66.43  62.14  43.78  57.49  30.83
OA  67.92  70.91  71.26  73.73  73.04  73.89  75.51  74.42  77.11
AA  62.84  44.13  67.69  63.50  67.32  63.56  67.78  65.96  67.93
Kappa  54.32  51.07  59.08  61.02  60.76  61.62  63.70  62.11  65.57
@@
Forest  90.88  93.78  97.06  97.38  97.47  90.04  97.07  96.52  97.55
Residential area  93.89  97.58  88.05  98.37  90.19  95.44  90.95  91.96  89.63
Industrial area  8.28  26.48  70.77  61.31  56.09  34.58  60.07  60.98  60.87
Low plants  91.97  97.67  96.88  92.63  97.38  90.68  95.57  96.64  97.94
Allotment  38.24  52.77  41.58  49.33  21.10  51.82  15.92  23.02  25.31
Commercial area  1.40  24.66  24.24  3.54  11.31  28.63  10.95  11.44  7.14
Water  10.82  47.51  81.82  26.61  87.59  17.65  84.64  83.92  88.09
OA  84.53  90.62  91.14  90.66  90.44  86.65  89.91  90.83  91.30
AA  47.92  62.92  62.51  61.30  62.91  58.40  61.27  62.36  64.42
Kappa  77.13  86.33  87.06  86.47  86.17  80.79  85.39  86.74  87.39
@@
Health grass  87.25  86.20  71.85  75.46  94.87  84.55  74.15  86.24  83.78
Stressed grass  96.62  93.34  99.25  95.39  93.89  89.13  96.94  93.91  94.56
Artificial turf  93.99  93.38  96.55  97.04  85.01  94.30  96.18  92.13  97.97
Evergreen trees  88.95  92.34  96.10  90.67  86.89  94.36  90.22  91.23  90.48
Deciduous trees  67.68  73.69  95.42  79.39  69.24  70.15  78.61  74.07  79.53
Bare earth  85.83  96.60  99.47  98.51  93.08  71.79  95.94  98.68  98.33
Water  98.26  99.12  61.57  93.07  98.43  90.82  97.92  66.74  95.27
Residential buildings  72.83  81.23  82.24  80.47  90.84  80.92  89.35  84.54  85.76
Non-residential buildings  99.71  99.29  99.56  99.72  99.39  99.68  99.14  99.42  99.47
Roads  76.15  85.40  82.39  89.61  85.87  88.06  82.86  85.43  89.96
Sidewalks  73.35  78.06  81.08  76.19  79.24  75.48  83.45  73.86  79.42
Crosswalks  7.70  22.68  13.56  22.99  18.02  19.58  19.36  14.49  26.85
Major thoroughfares  84.03  79.88  77.67  86.89  84.11  84.19  86.22  86.10  86.80
Highways  80.15  79.11  76.00  80.47  92.84  80.80  80.09  80.91  80.96
Railways  97.34  96.55  98.75  98.00  93.70  95.21  97.29  97.82  98.69
Paved parking lots  90.50  92.07  96.17  94.03  85.61  94.78  93.60  91.79  91.25
Unpaved parking lots  90.84  62.29  78.55  97.68  47.13  85.75  55.89  96.47  98.83
Cars  83.87  85.67  91.39  89.83  82.38  87.53  87.03  88.31  93.75
Trains  91.71  87.75  94.09  94.68  82.92  94.99  90.53  96.37  94.79
Stadium seats  91.69  95.76  95.48  76.02  98.36  86.86  97.63  89.51  99.58
OA  86.75  90.25  90.05  90.84  91.41  90.17  91.62  90.42  92.38
AA  92.70  94.25  93.13  95.34  93.67  94.17  95.10  94.69  95.51
Kappa  83.10  87.43  87.19  88.25  88.90  87.39  89.15  87.68  90.16

### v2

@@ III-C MSpa-Mamba Block
MSpa-SSM. As shown in Fig. 2(b), MSpa-SSM is the critical part of the MSpa-Mamba block. In MSpa-SSM, features at multiple scales are generated through DWConv with different stride values. These multi-scale feature maps are then handled by four scanning routes within SSM [59]. The scanning routes are divided into two groups: two maintain the original resolution and are processed by the SSM, while the others are downsampled, processed by the SSM, and then upsampled. This strategy reduces the overall feature volume by generating more compact feature representations, effectively preventing the accumulation of redundant information across all scanning routes. To be more specific, we use DWConv with strides of 1 and 2 to obtain feature map $\mathbf{Z}_{1}\in\mathbb{R}^{H\times W\times C}$ and $\mathbf{Z}_{2}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times C}$, respectively. Next, $\mathbf{Z}_{1}$ and $\mathbf{Z_{2}}$ are fed into the SSM as follows:
$[\mathbf{Y}_{1},\mathbf{Y}_{2}]=\textrm{SSM}(\sigma_{1}(\mathbf{Z}_{1}),\sigma_{2}(\mathbf{Z}_{1})),$ (9)
$[\mathbf{Y}_{3},\mathbf{Y}_{4}]=\textrm{SSM}(\sigma_{3}(\mathbf{Z}_{2}),\sigma_{4}(\mathbf{Z}_{2})),$ (10)
where $\sigma$ represents the transformation that reshapes the input features from $H\times W\times C$ to $HW\times C$, where $HW$ specifies the length of the spatial sequence to be processed by the SSM module. As illustrated in Fig. 2, the input is scanned in four distinct directions: the first scans rows first and then columns, while the second scans columns first and then rows. The third and fourth directions are the reverse of the first and second, respectively. Consequently, $\mathbf{Y}$ denotes the resulting processed sequence.
The obtained sequences are converted back into 2D feature maps, and the downsampled feature maps are interpolated for merging as follows:
@@
where $\beta$ is the inverse transformation of $\sigma$, reshaping the input data from $\mathbb{R}^{HW\times C}$ back to $\mathbb{R}^{H\times W\times C}$. Here, $\mathbf{Z}^{\prime}$ represents the feature map enhanced by the SSM module, and ‘Inter ’refers to the interpolation operation.
III-D Spe-Mamba Block
HSI covers a much larger spectral range with a higher spectral resolution. Effectively modeling the spectral relationship is important. In this paper, we design a Spe-Mamba block for spectral feature modeling.
@@ III-D Spe-Mamba Block
The Spe-Mamba shares a similar structure with the MSpa-Mamba, with the primary distinction being the SSM layer. In MSpa-SSM, features are transformed into 1D sequences along the spatial dimension, while in Spe-Mamba, they are converted into 1D sequences along the channel dimension, with scanning limited to two directions. Before being fed into Spe-SSM, the feature dimensions are reshaped from $\mathbb{R}^{H\times W\times C}$ to $\mathbb{R}^{C\times HW}$, where $C$ specifies the length of the sequence processed by the SSM module.
@@
III-E Fus-Mamba Block
Traditional SSMs typically handle single inputs and struggle to effectively integrate cross-modal information from multi-source data. To address this limitation, we propose the Fus-Mamba block, which integrates features from different modalities. As shown in Fig. 3, HSI and LiDAR/SAR features undergo a similar processing flow to that of MSpa-Mamba and Spe-Mamba. Multi-source features are then fed into the Fusion-SSM (Fus-SSM) block for cross-modal feature fusion and interaction.
The Fus-SSM serves as the core of Fus-Mamba, with its structure detailed in Fig. 3(b). Fus-SSM is designed symmetrically. In the top branch, the sequenced HSI feature $\mathbf{F}_{h}\in\mathbb{R}^{P\times C}$ generates projection and timescale parameters, while the sequenced LiDAR/SAR feature $\mathbf{F}_{x}\in\mathbb{R}^{P\times C}$ is processed to produce the output feature $\mathbf{F}_{xo}$. The bottom branch mirrors this structure: here, $\mathbf{F}_{x}$ generates the projection and timescale parameters, and $\mathbf{F}_{h}$ is processed to generate the output feature $\mathbf{F}_{ho}$.
Algorithm 1 outlines the computation process for $\mathbf{F}_{ho}$. Parameters A and D are initialized at the start, while B, C, and $\Delta$ are derived from linear layers applied to the sequence $\textbf{F}_{x}$ from another modality. In the original Mamba model, B, C, and $\Delta$ are derived from the sequence being processed, whereas in Fus-SSM, B, C, and $\Delta$ come from a different modality, specifically $\textbf{F}_{x}$. Since the computation for $\mathbf{F}_{xo}$ closely resembles that of $\mathbf{F}_{ho}$, it is omitted for brevity.
Comparison with Cross Mamba Block (CroMB): Although the Fus-Mamba Block shares conceptual similarities with the Cross Mamba Block (CroMB) in Sigma[65], it incorporates a more integrated cross-modal design. In CroMB, components $\mathbf{\overline{A}}$ and $\mathbf{\overline{B}}$ are generated solely from the sequence being processed, with only $\mathbf{C_{X}}$ relying on the sequence from the alternate modality. In contrast, our Fus-SSM utilizes features from the alternative modality to generate all three parameters ($\mathbf{\overline{A}}$, $\mathbf{\overline{B}}$ and $\mathbf{C}$) when processing a given sequence. This approach enables the Fus-Mamba Block to capture deeper interactions and dependencies between modalities, fully leveraging cross-modal information to improve feature alignment and enhance overall model performance.
@@ III-E Fus-Mamba Block
5: /* A and D are randomly initialized */
6:$\mathbf{B}:{\rm{{\color[rgb]{0,0.390625,0}(P,N)}}}\leftarrow\mathbf{Linear}(\mathbf{F}_{x})$
7:$\mathbf{C}:{\rm{{\color[rgb]{0,0.390625,0}(P,N)}}}\leftarrow\mathbf{Linear}(\mathbf{F}_{x})$
@@
12:$\mathbf{F}_{ho}\leftarrow{\rm{SSM}(\mathbf{\overline{A}},\mathbf{\overline{B}},\mathbf{C},\mathbf{D})}(\mathbf{F}_{h})$
@@ IV-A Dataset Description
To assess the effectiveness of our proposed MSFMamba, we applied it on four multi-source remote sensing datasets: Berlin, Augsburg, Houston 2018, and Houston 2013. The Berlin and Augsburg datasets are used for hyperspectral and SAR data classification, while the Houston 2013 and Houston 2018 datasets are used for hyperspectral and LiDAR data classification.
TABLE I: The number of training and test samples for the Berlin, Augsburg, Houston2018 and Houston2013 datasets.
@@ 20 Stadium seats 1000 26296
Houston2013 dataset
No. Name Color Train Test
1 Health grass 198 1053
2 Stressed grass 190 1064
3 Synthetic grass 192 505
4 Trees 188 1056
5 Soil 186 1056
6 Water 182 143
7 Residential 196 1072
8 Commercial 191 1053
9 Road 193 1059
10 Highway 191 1036
11 Railway 181 1054
12 Parking lot 1 192 1041
13 Parking lot 2 184 285
14 Tennis court 181 247
15 Running track 187 473
Total 2832 12197
Figure 4: Visualization of the four datasets used in our experiments.
Berlin Dataset. It covers the urban and rural areas of Berlin. The hyperspectral data is simulated Environmental Mapping and Analysis Program(EnMAP) data based on Hyperspectral Mapper(HyMap) data. The SAR data is Sentinel-1 dual-Pol (VV–VH) single look complex (SLC) product obtained from the European Space Agency (ESA) [66]. The HSI has 797$\times$220 pixels, 244 spectral bands in the wavelength range of 400–2500 nm [67]. The SAR image has 1723$\times$476 pixels. The nearest neighbor interpolation was employed to match the spatial resolution of the HSI and SAR data.
@@ 15 Running track 187 473
Houston2013 Dataset. The Houston2013 dataset, part of the 2013 IEEE Geoscience and Remote Sensing Society Data Fusion Contest, provides a unique perspective on urban land cover in Houston, Texas, and its surrounding areas. Captured at a spatial resolution of 2.5 meters, the dataset includes hyperspectral imagery with 144 spectral bands covering wavelengths from 380 to 1050 nm, along with LiDAR data that provides precise elevation information. Together, these data sources encompass 15 distinct land cover types, offering valuable insights for remote sensing and urban analysis.
Table LABEL:table_dataset enumerates the number of samples for both training and testing on the four datasets. Fig. 4 displays the HSI data in false color, along with the LiDAR/SAR images and the ground truth. The evaluation metrics used in this paper include Overall Accuracy(OA), Average Accuracy(AA), and Kappa. OA represents the proportion of correctly classified samples out of all the samples in the dataset, providing a measure of the model’s overall performance. AA is the average recall rate across all classes, reflecting the model’s effectiveness in handling different categories, particularly in imbalanced datasets. Kappa, on the other hand, is a statistical measure that quantifies the agreement between the predicted and true labels, adjusted for random chance, offering a more robust evaluation of classifier performance than simple accuracy, especially when dealing with imbalanced or noisy data.
@@ IV-B Parameter Analysis
Number of Spatial-Spectral Mamba Module. The spatial-spectral Mamba module is the critical part of our MSFMamba, and the number of modules, $L$, is a key parameter that affects classification performance. We tested the optimal number of spatial-spectral Mamba modules for each dataset. Fig. 5 illustrates the relationship between the OA value and $L$ on each dataset. Our MSFMamba achieves the best classification performance on the Augsburg and Houston2013 datasets when $L=1$. For the Berlin and Houston2018 datasets, the best performance is observed when $L=2$. Therefore, in our subsequent experiments, we set $L=1$for the Augsburg and Houston2013 datasets and $L=2$ for the Berlin and Houston2018 datasets.
Number of Principal Components in HSI. We use PCA to reduce the spectral redundancy of the input HSI, with the number of principal components,$N_{p}$, being a critical parameter. We tested different values of $N_{p}$ ranging from 20 to 40, and the experimental results are shown in Fig. 6. It can be seen that when $N_{p}=30$, the OA values for the Augsburg and Houston2018 datasets reach their highest. For the Berlin and Houston2013 datasets, the best OA values are achieved when $N_{p}=35$.
Number of States in SSM. The number of states is an important parameter in the state space model. We tested five different values for the number of states, ranging from 4 to 20. As shown in Fig. 7, the optimal number of states for the Houston2018, Houston2013, and Berlin datasets is 16, while for the Augsburg dataset, the optimal number of states is 8.
The best settings of these experiments were used to evaluate our MSFMamba for comparison with other state-of-the-art methods. All experiments were performed with an RTX 4090 GPU and 32GB of RAM using PyTorch.
@@
Forest  81.75  86.24  83.27  82.04  88.35  84.71  83.18  82.21  82.39
Residential area  76.26  91.38  72.07  77.78  74.44  76.19  78.80  78.16  82.05
Industrial area  39.67  19.76  46.66  47.96  48.21  44.15  43.59  47.79  48.91
Low plants  49.78  20.00  72.08  77.78  71.41  77.96  76.48  78.04  79.72
Soil  89.42  48.72  77.94  87.85  80.75  67.13  82.92  77.26  85.22
Allotment  54.36  38.89  70.62  44.75  48.35  61.95  56.46  66.07  71.25
Commercial area  4.65  18.47  36.48  29.85  27.97  29.28  18.96  28.90  31.08
Water  41.93  29.61  54.64  60.09  75.16  56.98  44.16  50.22  38.48
OA  67.60  70.91  70.38  73.69  71.65  72.59  73.42  73.98  76.92
AA  54.72  44.13  64.22  63.51  64.33  62.39  60.57  63.58  64.88
Kappa  50.96  51.07  57.73  61.02  58.74  59.72  60.29  61.19  64.88
Figure 8: Classification results of different methods for the Berlin dataset.
@@
Forest  90.88  94.34  98.10  97.38  97.56  90.04  97.26  97.84  97.17
Residential area  93.89  92.56  99.08  98.37  99.16  95.44  98.63  98.73  98.15
Industrial area  8.28  47.70  12.19  61.31  61.38  34.58  36.61  43.94  50.26
Low plants  91.97  85.96  91.78  92.63  83.47  90.68  93.54  94.11  95.52
Allotment  38.24  49.33  45.12  49.33  42.64  51.82  51.43  52.39  53.35
Commercial area  1.40  11.52  1.22  3.54  5.58  28.63  7.97  6.80  2.63
Water  10.82  45.27  24.09  26.61  45.81  17.65  46.34  48.54  49.97
OA  84.53  85.31  88.22  90.66  88.25  86.65  90.33  91.06  91.38
AA  47.93  62.93  53.08  61.31  62.23  58.41  61.68  63.19  63.31
Kappa  77.13  86.33  86.47  86.47  83.13  80.79  86.03  87.06  87.45
Figure 9: Classification results of different methods for the Augsburg dataset.
@@
Health grass 94.84 60.69 99.15 94.29 96.9 93.34 97.38 89.53 90.38
Stressed grass 92.6 93.85 84.35 92.71 91.82 93.44 89.4 93.5 94.04
Artificial turf 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
Evergreen trees 98.8 99.02 98.13 98.92 99.29 95.87 98.9 99.1 99.34
Deciduous trees 97.12 94.73 86.39 97.77 99.52 96.85 98.89 98.49 99.11
Bare earth 99.61 99.87 99.33 100.0 99.98 100.0 100.0 99.96 100.0
Water 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
Residential buildings 93.21 95.51 97.17 97.36 95.16 97.39 95.3 95.31 96.27
Non-residential buildings 91.3 93.64 93.93 93.29 95.02 93 95.25 93.86 95.58
Roads 61.06 70.94 70.17 75.73 72.72 72.18 72.77 72.11 75.9
Sidewalks 75.91 79.11 72.11 83.67 77.18 69.69 73.27 80.69 81.43
Crosswalks 85.31 86.06 82.47 94.23 93.14 88.12 98.95 97.69 96.22
Major thoroughfares 72.77 76.12 89.44 81.2 84.67 86.12 80.9 79.05 86.89
Highways 95.86 98.89 98.53 98.91 99.53 99.5 98.01 96.86 98.63
Railways 99.78 99.88 99.21 99.94 99.90 99.87 99.74 99.9 99.69
Paved parking lots 90.74 95.88 95.62 98.37 99.18 97.02 97.07 97.74 98.94
Unpaved parking lots 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
Cars 98.47 94.25 96.77 99.09 96.64 98.13 99.01 98.74 97.81
Trains 99.9 99.84 100.0 99.46 100.0 99.98 99.97 99.74 99.97
Stadium seats 99.92 100.0 100.0 99.98 99.89 99.99 99.98 99.99 100.0
OA 86.95 89.09 90.05 91.02 91.24 89.98 90.56 90.31 92.38
AA 92.36 91.91 93.14 95.24 95.02 94.02 94.74 94.61 95.51
Kappa 83.39 85.96 87.2 88.46 88.71 87.14 87.81 87.55 90.16
Figure 10: Classification results of different methods for the Houston 2018 dataset.
TABLE V: Classification performance of different methods on the Houston 2013 dataset.
Class TBCNN FusAtNet $S^{2}$et DFINet AsyFFNet ExViT HCT MACN MSFMamba
Health grass 82.72 82.53 81.48 82.91 82.72 82.72 81.1 82.90 82.72
Stressed grass 83.18 85.15 84.68 85.15 83.93 85.15 86.09 100.0 99.44
Synthetic grass 100.0 98.81 99.01 99.60 100.0 99.21 99.80 99.80 98.81
Trees 94.03 92.90 92.14 92.33 91.95 91.95 98.01 98.11 98.39
Soil 99.15 100.0 100.0 99.81 100.0 100.0 100.0 100.0 100.0
Water 99.30 98.60 100.0 100.0 95.80 98.60 95.10 95.80 95.80
Residential 79.66 85.45 89.93 93.47 96.18 86.66 83.49 78.17 82.18
Commercial 55.27 81.81 90.18 82.58 81.14 91.05 86.43 80.46 94.23
Road 75.83 83.76 92.63 89.05 86.12 91.12 86.69 90.93 93.96
Highway 62.55 53.67 64.86 56.56 64.58 64.86 79.05 55.88 79.92
Railway 96.87 79.65 97.60 94.15 86.76 77.06 95.97 99.33 93.47
Parking lot 1 86.55 91.26 88.09 94.72 90.59 88.76 98.46 99.71 96.73
Parking lot 2 53.68 81.05 91.93 89.12 92.28 87.72 90.88 88.77 90.88
Tennis court 98.79 100.0 100.0 100.0 95.14 97.17 100.0 100.0 100.0
Running track 98.10 98.73 100.0 100.0 100.0 98.52 100.0 100.0 100.0
OA 82.91 85.32 89.56 88.59 87.95 87.42 90.66 89.81 92.86
AA 84.38 87.56 91.50 90.63 89.81 89.37 92.07 91.33 93.77
Kappa 81.43 84.12 88.69 87.64 86.96 86.38 89.86 88.93 92.25
Figure 11: Classification results of different methods for the Houston 2013 dataset.